# Orbital-Selective Diffuse Magnetic Fluctuations in Sr2RuO4: a Unified Theoretical Picture

### v1

@@ Abstract
The quasi-two-dimensional material Sr2RuO4 has been the focus of extensive experimental and theoretical research, as it is a paradigmatic example of a correlated system that exhibits unconventional superconductivity and intriguing magnetic properties. The interplay between these two effects has sparked significant debates, especially on the strength of the spin excitations. We show that self-consistently incorporating spatial magnetic fluctuations into our theoretical framework significantly reduces the many-body effects in the system. Consistent with experimental observations, this reduction destabilizes the magnetic ordering in Sr2RuO4, which is not found in our calculations in contrast to previous theoretical studies. This resolution of the long-standing discrepancy between theory and experiment is supported by a theoretical calculation of the spin susceptibility that closely matches the experimental results.
@@
In strongly correlated systems, typically containing partially filled or electron shells, Coulomb interactions between the electrons play a dominant role in determining the materials properties. The behavior of electrons cannot be described independently from one another, as is mostly the case for conventional metals, and complex phenomena emerge, including high-temperature superconductivity, Mott insulating states, or exotic magnetic phases.
The layered ruthenate Sr2RuO4 has become a drosophila compound in the field of correlated systems, due to its simple fabrication in large single crystals, forming notably clean samples. These attributes facilitate accurate experimental measurements and characterisation of the material, which has thus attracted significant scientific attention. Furthermore, the quasi-two-dimensional nature of Sr2RuO4 has established it as a benchmark material for experimental research.
The material’s Fermi surface (FS), shown in Fig. 1, consists of three Fermi sheets, originating from the three orbitals: the quasi-one-dimensional and , and the quasi-two-dimensional Mackenzie et al. (1996); Bergemann et al. (2000); Damascelli et al. (2000); Tamai et al. (2019). Angle-resolved photoemission spectroscopy (ARPES) and quantum oscillations measurements reveal that Sr2RuO4 is a correlated metal Ingle et al. (2005); Damascelli et al. (2000); Mackenzie et al. (1996). Although a significant redistribution of the spectral weight and an enhancement of it around the Fermi energy are observed Lu et al. (1996); Yokoya et al. (1996), the band structure and FS remain relatively close Mazin and Singh (1997); Pavarini and Mazin (2006) to those predicted by density functional theory (DFT) Kohn and Sham (1965). From the theoretical perspective, the compound is considered to be a paradigmatic correlated Fermi liquid, with strong signatures of Hund’s metal physics Mravlje et al. (2011); Georges et al. (2013); Kugler et al. (2020); Hunter et al. (2023); Blesio, Germán and Beck, Sophie and Gingras, Olivier and Georges, Antoine and Mravlje, Jernej (2024); Suzuki, H and Wang, L and Bertinshaw, J and Strand, Hugo UR and Käser, S and Krautloher, M and Yang, Z and Wentzell, N and Parcollet, O and Jerzembeck, F and others (2023). Therefore, a number of studies have been conducted using this material as a testbed for realistic many-body calculations.
An early electronic structure calculation using DFT combined with the state-of-the-art dynamical mean field theory (DMFT) Georges et al. (1996), that accounts for only local correlation effects, has nearly reproduced the experimental spectral function Liebsch and Lichtenstein (2000). Through DMFT studies, many important advancements have been made in understanding the physics of Sr2RuO4. The FS has been reproduced reasonably well, along with the mass enhancement for each orbital Mravlje et al. (2011); Hunter et al. (2023); Zhang et al. (2016). Furthermore, recent studies Tamai et al. (2019); Hunter et al. (2023) showed that the self-energy, derived from photoemission spectra measurements at the k-points of the Brillouin Zone (BZ) corresponding to the FS, is predominantly local for the orbitals. In contrast, the orbital exhibits some momentum dependence, which becomes apparent at frequencies above meV. These findings suggest that, overall, the compound’s single-particle properties are not particularly remarkable and are adequately captured by DMFT.
It is therefore very surprising that the material’s two-particle quantities are, in contrast, not well understood. Experimental measurements have shown that at low temperatures (below K), Sr2RuO4 exhibits superconductivity of an unconventional nature Maeno et al. (1994); Mackenzie and Maeno (2003); Maeno et al. (2012), while recently an increase of has been investigated when the system is under uniaxial strain H. M. L. Noad and K. Ishida and Y.-S. Li and E. Gati and V. Stangier and N. Kikugawa and D. A. Sokolov and M. Nicklas and B. Kim and I. I. Mazin and M. Garst and J. Schmalian and A. P. Mackenzie and C. W. Hicks (2023); Maeno et al. (2024); Clifford W. Hicks and Daniel O. Brodsky and Edward A. Yelland and Alexandra S. Gibbs and Jan A. N. Bruin and Mark E. Barber and Stephen D. Edkins and Keigo Nishimura and Shingo Yonezawa and Yoshiteru Maeno and Andrew P. Mackenzie (2014); You-Sheng Li and Naoki Kikugawa and Dmitry A. Sokolov and Fabian Jerzembeck and Alexandra S. Gibbs and Yoshiteru Maeno and Clifford W. Hicks and Jörg Schmalian and Michael Nicklas and Andrew P. Mackenzie (2021). At higher temperatures, significant magnetic fluctuations are found, and they are believed to be the source of the superconducting pairing mechanism in Sr2RuO4, making their accurate description essential. Initially, the proximity of the compound to SrRuO3, which orders ferromagnetically Kanbayasi (1976); Maeno et al. (1994), suggested that the system forms a triplet superconducting state Duffy et al. (2000); Deguchi et al. (2004); Nelson et al. (2004). However, once it was discovered that the leading spin excitations correspond to an incommensurate antiferromagnetic wave vector Sidis et al. (1999); Braden, M. and Sidis, Y. and Bourges, P. and Pfeuty, P. and Kulda, J. and Mao, Z. and Maeno, Y. (2002); Braden et al. (2004); Steffens et al. (2019); Jenni et al. (2021), the scenario of a singlet state became prominent. Nevertheless, the situation remains unresolved, since conflicting experimental evidences suggest broken time-reversal symmetry Xia et al. (2006); Luke et al. (1998) and possibly a two-component superconducting state Maeno et al. (2012); Kivelson et al. (2020); Maeno et al. (2024). A number of theoretical works have been performed, aiming to solve the mystery of the superconducting state in Sr2RuO4 Rømer et al. (2019, 2022); Profe et al. (2024) but no definite conclusion has been drawn so far.
As seen in inelastic neutron scattering (INS) measurements, the leading spin excitations in Sr2RuO4 correspond to the incommensurate wave vector Sidis et al. (1999); Braden, M. and Sidis, Y. and Bourges, P. and Pfeuty, P. and Kulda, J. and Mao, Z. and Maeno, Y. (2002); Braden et al. (2004); Steffens et al. (2019); Jenni et al. (2021), which is directly linked to nesting vectors (Fig. 1 d) that connect different parts of the FS within the same and orbitals. These strong magnetic fluctuations imply the proximity to a spin density wave (SDW) instability. However, different experimental studies confirm the absence of magnetic ordering in the parent compound. Nevertheless, the presence of even small concentrations of impurities have been found to induce a magnetic transition Minakata and Maeno (2001); Braden et al. (2002); Kunkemöller, S. and Nugroho, A. A. and Sidis, Y. and Braden, M. (2014).
The behavior of two-particle quantities in Sr2RuO4, particularly spin excitations, has sparked debates among theorists. Within DFT calculations, the FS nesting arising from the compound’s crystal structure leads to an ordered SDW state with the incommensurate ordering vectors Mazin and Singh (1999). DMFT calculations of the spin susceptibility, incorporating local vertex corrections, have also reproduced the peaks at the incommensurate wave vectors Boehnke et al. (2018); Acharya et al. (2019); Strand et al. (2019). However, DMFT also predicts a transition to a SDW ordered state at a finite temperature K, in contrast to experimental findings Strand et al. (2019). Within DMFT, the transition appears to be suppressed only by introducing a relatively large spin-orbit coupling Strand et al. (2019). This result is somewhat inconclusive, as it is based on calculations conducted at relatively high temperatures. Therefore, the DMFT approximation seems to significantly overestimate the strength of magnetic fluctuations, which, according to experiments, should be less pronounced.
The key contribution of DMFT to calculating spin susceptibility is its ability to reveal the suppression of the magnetic signal at the edges of the BZ, particularly at the point, in agreement with experiments. Within DMFT, spin excitations, aside from the SDW peaks, are found to be quasi-local, or nearly constant in momentum space. However, this result does not fully align with experimental findings, where the magnetic response, in addition to the SDW peaks, exhibits a relatively broad dome centered at the point Steffens et al. (2019); Jenni et al. (2021, 2021), rather than a quasi-local background signal. The final missing component in these DMFT calculations, which could potentially reconcile the discrepancies between theory and experiment, is the self-consistent incorporation of spatial magnetic fluctuations, that are exceptionally strong in Sr2RuO4, into the theory. This effect has not been addressed yet due to the lack of an appropriate method.
@@
In this Letter, we resolve this long-standing problem by applying a recently developed advanced many-body approach that enables a self-consistent treatment of non-local collective electronic fluctuations beyond DMFT, within a realistic multi-orbital framework. We demonstrate that a self-consistent treatment of spatial spin excitations suppresses their strength and reduces the electronic correlations in Sr2RuO4, ultimately eliminating the magnetic ordering predicted by DMFT. Moreover, our results accurately reproduce the form of the spin susceptibility as deduced from experimental INS measurements.
@@
We study an effective three-band model (see the Supplemental Material (SM) SM ) corresponding to maximally localized orbitals derived from DFT 111 The DFT calculations were performed within the projected augmented wave formalism Blöchl (1994); Kresse and Joubert (1999) as implemented in the Vienna ab initio simulation package (VASP) Kresse and Furthmüller (1996, 1996). The exchange-correlation effects were considered within the generalized-gradient approximation functional in the PBE parametrization Perdew et al. (1996). We used the standard pseudopotentials that include 10, 16, and 6 valence electrons for Sr, Ru, and O, respectively. An energy cut-off of 400 eV for the plane-waves and a convergence threshold of eV were used in the calculations. The Wannier functions and the effective tight-binding Hamiltonian were constructed within the scheme of maximal localization Marzari and Vanderbilt (1997); Marzari et al. (2012) using the wannier90 package Mostofi et al. (2008). . We account for the on-site electronic interaction that is parametrized in the Kanamori form Kanamori (1963). The value of the intra-orbital Coulomb repulsion, eV, is chosen based on the constrained random-phase approximation (cRPA) analysis conducted in Vaugier et al. (2012). A Hund’s exchange coupling , crucial for Sr2RuO4, is selected by evaluating results across different values of . Specifically, we estimate the mass enhancement and spin susceptibility obtained for each , as detailed in the SM SM . We find that the correct mass enhancement is achieved for eV, and the accurate spin susceptibility for eV, leading to an optimal choice of eV. We do not include spin-orbit coupling (SOC) in our calculations because it is prohibitively expensive computationally. This approximation is supported by experimental evidences, suggesting that SOC effects become less significant at higher temperatures, due to thermal effects. This is seen by studying the corresponding anisotropy of the spin susceptibility, related to SOC, which diminishes within the temperature range to which we are restricted Eremin et al. (2002); Braden et al. (2004).
@@
In order to obtain a reliable description of the feedback of electronic correlations on the spectral and magnetic properties of the compound, a self-consistent method is required. We employ the dual triply irreducible local expansion (D-TRILEX) approach Stepanov et al. (2019); Harkov et al. (2021); Vandelli et al. (2022), which is a diagrammatic extension of DMFT Rohringer et al. (2018); Lyakhova et al. (2023). The decisive advantage of this method over other DMFT extensions is its computational efficiency, which enables calculations for multi-orbital systems Stepanov et al. (2021); Vandelli et al. (2023); Stepanov (2022); Stepanov and Biermann (2024); Stepanov et al. (2024a), such as Sr2RuO4. In the D-TRILEX framework, local electronic correlations are treated non-perturbatively through an exactly solved DMFT impurity problem Georges et al. (1996). The latter further serves as a reference system for the diagrammatic expansion that self-consistently incorporates non-local electronic correlations, including magnetic fluctuations. This approach allows the mutual influence of collective electronic fluctuations on single-particle quantities and vice versa Stepanov et al. (2022); Vandelli et al. (2024); Chatzieleftheriou et al. (2024); Stepanov et al. (2024b), yielding reliable results for both single- and two-particle observables in a broad regime of system’s parameters Harkov et al. (2021); Vandelli et al. (2022); Vandelli (2022).
@@
The second important outcome of our results is related to the overall behavior of the spin susceptibility across the BZ. As discussed above, according to experimental measurements the magnetic signal can be decomposed into the sum of the SDW peaks and a broad dome structure centered around the point, while DMFT calculations instead find a quasi-constant background signal besides the peaks. The D-TRILEX calculations reveal a significantly diminished spin susceptibility at the edges of the BZ, with a “cross”-like structure in momentum space of higher intensity appearing at the center of the zone, visible in the inset of Fig. 2. The overall structure of the susceptibility agrees very well with the experimental results Steffens et al. (2019). While early-unpolarized INS studies reported features of the spin excitation spectrum around Braden, M. and Sidis, Y. and Bourges, P. and Pfeuty, P. and Kulda, J. and Mao, Z. and Maeno, Y. (2002), the most recent polarized INS data Steffens et al. (2019); Jenni et al. (2021) unambiguously confirmed the existence of such fluctuations, but cannot resolve the actual structure around . Nevertheless, strong spin excitations around are found in metamagnetic Ca2-xSrxRuO4 Friedt, O. and Steffens, P. and Braden, M. and Sidis, Y. and Nakatsuji, S. and Maeno, Y. (2004); Steffens, P. and Sidis, Y. and Link, P. and Schmalzl, K. and Nakatsuji, S. and Maeno, Y. and Braden, M. (2007); Steffens, P. and Friedt, O. and Sidis, Y. and Link, P. and Kulda, J. and Schmalzl, K. and Nakatsuji, S. and Braden, M. (2011), where the isostructural Ca/Sr substitution is, however, accompanied by a rotation and deformation of the RuO6 octahedra Friedt, O. and Braden, M. and André, G. and Adelmann, P. and Nakatsuji, S. and Maeno, Y. (2001); Kriener, M. and Steffens, P. and Baier, J. and Schumann, O. and Zabel, T. and Lorenz, T. and Friedt, O. and Müller, R. and Gukasov, A. and Radaelli, P. G. and Reutler, P. and Revcolevschi, A. and Nakatsuji, S. and Maeno, Y. and Braden, M. (2005).
@@
In Fig. 3 we show the orbitally resolved static spin susceptibility , where . The results are calculated using D-TRILEX (blue) and DMFT (green) and plotted along the X--X path. Although the incommensurate SDW peaks are present in all orbital components due to mixing of different of orbital contributions to the susceptibility, the value of is the largest for the intra-orbital - and - components (a). This is expected, since the SDW peaks originate from the nesting of the FS of these 1D-like orbitals. Interestingly, the background dome-shaped magnetic signal, found experimentally and reproduced by D-TRILEX, appears to originate from the magnetic fluctuations within the 2D-like orbital (b). Instead, the DMFT results do not exhibit any significant signal around the point, and for all the components a quasi-local response is obtained besides the peaks. This result may have significant implications for the symmetry of the superconducting order parameter, because the magnetic fluctuations within the orbital are believed to be a key ingredient for the electron pairing Rømer et al. (2019); Profe et al. (2024). In particular, the presence of spin excitations at more than one wave-vector allows for a complex (possibly two-component) order parameter of the superconducting state Ghosh et al. (2021); Benhabib et al. (2021). Therefore, the spin singlet state Rømer et al. (2022); Moon (2023); Profe et al. (2024); Kaba and Sénéchal (2019); Ramires and Sigrist (2019) promoted by the incommensurate fluctuations at the wave vector might compete with a spin triplet state Rømer et al. (2019, 2022); Zhang et al. (2018); Gingras et al. (2019) related to the momenta around the center of the BZ.
@@
By analyzing the electron self-energy we find that is has a surprisingly small momentum-dependence along the FS, as discussed in the SM SM . Along the FS the / orbitals appear to have a practically constant self-energy, while the orbital exhibits a finite but small momentum-modulation. This result is in a qualitative agreement with the conclusion of a recent ARPES experiment Tamai et al. (2019). It is remarkable that the self-energy in Sr2RuO4 is not local, as it exhibits a noticeable momentum dependence perpendicularly to the FS.
@@
These excitations are found to be significantly strong in this material, such that DMFT calculations predict an ordered SDW state that is not observed experimentally.
The systematic incorporation of non-local correlations resolves this discrepancy between theory and experiments.
We demonstrate, that the self-consistent inclusion of magnetic fluctuations suppresses their strength by reducing the many-body effects in the system, and through that no magnetic ordering is realized in Sr2RuO4.
The overall behavior of the spin susceptibility in momentum space, deduced from the INS measurements, is well reproduced by our calculations.
@@
We speculate that the predicted momentum-space structure of the spin susceptibility favors the realization of a superconducting state with a mixed-component order parameter, in line with the most recent experimental evidence.

### v2

@@ Abstract
The quasi-two-dimensional material Sr2RuO4 is a paradigmatic example of a correlated system that exhibits unconventional superconductivity and intriguing magnetic properties. The interplay between these two effects and the resulting strength and nature of spin fluctuations and their role for the properties of the compound have sparked significant debates. Here, elaborating a theory that self-consistently incorporates spatial magnetic fluctuations into a realistic many-body description, we show that these fluctuations significantly reduce many-body correlations in the system, thereby preventing magnetic ordering in Sr2RuO4, in agreement with experimental observations. Our conclusion is supported by a theoretical calculation of the spin susceptibility that closely matches the experimental results. We obtain finite peaks at the incommensurate wave vectors, a broad dome-shaped structure centered around the point and a diminished magnetic response at the edges of the BZ. We identify the orbital character of the unusual dome structure as resulting predominantly from the 2D-like -orbital, which is believed to be responsible for the superconductivity.
@@
In materials with strong electronic correlations, typically containing partially filled electronic - or -shells, Coulomb interactions between the electrons play a dominant role in determining their properties. The behavior of electrons cannot be described independently from one another, as is mostly the case for conventional metals. Instead, complex collective phenomena emerge, including high-temperature superconductivity, Mott insulating states, or exotic magnetic phases.
The layered ruthenate Sr2RuO4 has attracted significant attention due to its structural similarity to high-temperature superconducting cuprates Keimer et al. (2015) and its unconventional superconductivity below K Maeno et al. (1994); Mackenzie and Maeno (2003); Maeno et al. (2012). Currently, ongoing efforts focus on understanding the enhancement of under uniaxial strain H. M. L. Noad and K. Ishida and Y.-S. Li and E. Gati and V. Stangier and N. Kikugawa and D. A. Sokolov and M. Nicklas and B. Kim and I. I. Mazin and M. Garst and J. Schmalian and A. P. Mackenzie and C. W. Hicks (2023); Maeno et al. (2024); Clifford W. Hicks and Daniel O. Brodsky and Edward A. Yelland and Alexandra S. Gibbs and Jan A. N. Bruin and Mark E. Barber and Stephen D. Edkins and Keigo Nishimura and Shingo Yonezawa and Yoshiteru Maeno and Andrew P. Mackenzie (2014); You-Sheng Li and Naoki Kikugawa and Dmitry A. Sokolov and Fabian Jerzembeck and Alexandra S. Gibbs and Yoshiteru Maeno and Clifford W. Hicks and Jörg Schmalian and Michael Nicklas and Andrew P. Mackenzie (2021). Despite extensive research, however, the precise nature of the superconducting state in Sr2RuO4 remains unresolved. Conflicting experimental evidence suggests scenarios ranging from a single Chronister et al. (2021); Li et al. (2022); Jerzembeck et al. (2022); Palle et al. (2023) to broken time-reversal symmetry Xia et al. (2006); Luke et al. (1998), and a two-component Maeno et al. (2012); Kivelson et al. (2020); Maeno et al. (2024) order parameter. Numerous theoretical studies have tried to resolve this issue Rømer et al. (2019); Käser et al. (2022); Rømer et al. (2022); Profe et al. (2024), but no consensus has been reached so far.
At higher temperatures, significant magnetic fluctuations are found, and they are believed to be the source of the superconducting pairing mechanism in Sr2RuO4, calling urgently for their accurate theoretical description. As seen in inelastic neutron scattering (INS) measurements, the leading spin excitations in Sr2RuO4 correspond to the incommensurate wave vector Sidis et al. (1999); Braden et al. (2002a, 2004); Steffens et al. (2019); Jenni et al. (2021). Remarkably, although spin excitations in this material are exceptionally strong, different experimental studies confirm the absence of magnetic ordering in the parent compound. Nevertheless, Sr2RuO4 resides in close proximity to an ordered state, which can be induced already by a small concentration of impurities Minakata and Maeno (2001); Braden et al. (2002b); Kunkemöller et al. (2014), or by applying strain Grinenko et al. (2021); Li et al. (2022).
The description of spin excitations in Sr2RuO4 has sparked debates among theorists. This material is a paradigmatic correlated metal that requires an accurate non-perturbative treatment of electronic interactions. Conventional theoretical methods correctly capture the leading spin fluctuations associated with the incommensurate wave vector Mazin and Singh (1999); Boehnke et al. (2018); Acharya et al. (2019); Strand et al. (2019). However, they significantly overestimate the strength of magnetic fluctuations, predicting a transition to a magnetically ordered state that is not realized in nature. The final missing component in state-of-the-art theoretical calculations, which could potentially reconcile the discrepancies between theory and experiment, is a self-consistent incorporation of spatial magnetic fluctuations into the theory. This effect has not been addressed yet due to the lack of appropriate theoretical tools.
@@
In this Letter, we resolve this long-standing problem by applying a recently developed advanced many-body approach that enables a self-consistent treatment of non-local collective electronic fluctuations within a realistic multi-orbital framework. We demonstrate that a self-consistent treatment of spatial spin excitations suppresses their strength and reduces the electronic correlations in Sr2RuO4, ultimately eliminating the magnetic ordering predicted by state-of-the-art approaches. Moreover, our results accurately reproduce the form of the spin susceptibility as deduced from experimental INS measurements, allowing us to detect an intriguing orbital dependence of the latter.
The Fermi surface (FS) of Sr2RuO4, shown in Fig. 1, consists of three Fermi sheets, originating from the three orbitals: the quasi-one-dimensional (1D) and , and the quasi-two-dimensional (2D) Mackenzie et al. (1996); Bergemann et al. (2000); Damascelli et al. (2000); Tamai et al. (2019). Angle-resolved photoemission spectroscopy (ARPES) and quantum oscillations measurements reveal that Sr2RuO4 is a correlated metal Ingle et al. (2005); Damascelli et al. (2000); Mackenzie et al. (1996). Although a significant redistribution of the spectral weight and an enhancement of it around the Fermi energy are observed Lu et al. (1996); Yokoya et al. (1996), the band structure and FS remain relatively close Mazin and Singh (1997); Pavarini and Mazin (2006) to those predicted by density functional theory (DFT) Kohn and Sham (1965). From the theoretical perspective, the compound is considered to be a paradigmatic correlated Fermi liquid, with strong signatures of Hund’s metal physics Mravlje et al. (2011); Georges et al. (2013); Kugler et al. (2020); Hunter et al. (2023); Blesio, Germán and Beck, Sophie and Gingras, Olivier and Georges, Antoine and Mravlje, Jernej (2024); Suzuki et al. (2023). Therefore, a number of studies have been conducted using this material as a testbed for realistic many-body calculations.
Early electronic structure calculations using DFT combined with dynamical mean-field theory (DMFT) Georges et al. (1996), that accounts for local correlation effects, have nearly succeeded to reproduce the experimental spectral function Liebsch and Lichtenstein (2000). Since then, several DMFT studies have allowed for many important advances in our understanding of the physics of Sr2RuO4 Mravlje et al. (2011); Hunter et al. (2023); Zhang et al. (2016). Recent studies Tamai et al. (2019); Hunter et al. (2023) showed that the self-energy, derived from photoemission measurements at the k-points of the Brillouin Zone (BZ) corresponding to the FS, is predominantly local for the orbitals. In contrast, the orbital exhibits some momentum dependence, which becomes apparent at frequencies above meV Tamai et al. (2019). These findings suggest that, overall, the compound’s single-particle properties are not particularly remarkable and are adequately captured by DMFT.
It is therefore very surprising that the material’s two-particle quantities, particularly spin excitations, are in contrast not well understood. Within DFT calculations, the FS nesting arising from the compound’s crystal structure leads to an ordered spin-density wave (SDW) state with the incommensurate wave vectors Mazin and Singh (1999). DMFT calculations of the spin susceptibility, incorporating local vertex corrections, have also reproduced the peaks at the vectors Boehnke et al. (2018); Acharya et al. (2019); Strand et al. (2019). However, DMFT strongly overestimates the strength of magnetic fluctuations, compared to experiments, predicting a transition to a SDW ordered state at a finite temperature K Strand et al. (2019). While it has been shown that spin-orbit coupling (SOC) can suppress this magnetic transition in DMFT, achieving this requires a rather large coupling strength Strand et al. (2019). Additionally, the SOC is not expected to significantly influence the spin susceptibility at the relatively high temperatures at which the DMFT calculations have been conducted Eremin et al. (2002); Braden et al. (2004).
The key contribution of DMFT to calculating the spin susceptibility is its ability to reveal the suppression of the magnetic signal at the edges of the BZ, particularly at the point, in agreement with experiments. Within DMFT, spin excitations, aside from the peaks, are found to be quasi-local, or nearly constant in momentum space. However, this result does not fully align with experimental findings, where the magnetic response, in addition to the peaks, exhibits a relatively broad dome centered at the point Steffens et al. (2019); Jenni et al. (2021, 2021), rather than a quasi-local background signal.
@@
To accurately describe magnetic fluctuations in Sr2RuO4, we consider an effective three-band model (see the Supplemental Material (SM) SM ) corresponding to maximally localized orbitals derived from DFT 111 The DFT calculations were performed within the projected augmented wave formalism Blöchl (1994); Kresse and Joubert (1999) as implemented in the Vienna ab initio simulation package (VASP) Kresse and Furthmüller (1996, 1996). The exchange-correlation effects were considered within the generalized-gradient approximation functional in the PBE parametrization Perdew et al. (1996). We used the standard pseudopotentials that include 10, 16, and 6 valence electrons for Sr, Ru, and O, respectively. An energy cut-off of 400 eV for the plane-waves and a convergence threshold of eV were used in the calculations. The Wannier functions and the effective tight-binding Hamiltonian were constructed within the scheme of maximal localization Marzari and Vanderbilt (1997); Marzari et al. (2012) using the wannier90 package Mostofi et al. (2008). . We account for the on-site electronic interaction that is parametrized in the Kanamori form Kanamori (1963). The value of the intra-orbital Coulomb repulsion, eV, is chosen based on the constrained random-phase approximation (cRPA) analysis conducted in Vaugier et al. (2012). A Hund’s exchange coupling , crucial for Sr2RuO4, is selected by evaluating results across different values of . Specifically, we estimate the mass enhancement and spin susceptibility obtained for each , as detailed in the SM SM . We find that the correct mass enhancement is achieved for eV, and the accurate spin susceptibility for eV, leading to an optimal choice of eV. We do not include spin-orbit coupling (SOC) in our calculations because it is prohibitively expensive computationally. This approximation is supported by experimental evidences, suggesting that SOC effects become less significant at higher temperatures, due to thermal effects.
@@
In order to obtain a reliable description of the feedback of electronic correlations on the spectral and magnetic properties of the compound, a self-consistent method is required. We employ the dual triply irreducible local expansion (D-TRILEX) approach Stepanov et al. (2019); Harkov et al. (2021); Vandelli et al. (2022), which is a diagrammatic extension of DMFT Rohringer et al. (2018); Lyakhova et al. (2023). The decisive advantage of this method over other DMFT extensions is its computational efficiency, which enables calculations for multi-orbital systems Stepanov et al. (2021); Vandelli et al. (2023); Stepanov (2022); Stepanov and Biermann (2024); Stepanov et al. (2024a), such as Sr2RuO4. In the D-TRILEX framework, local electronic correlations are treated non-perturbatively via a DMFT impurity problem Georges et al. (1996), which is solved using the w2dynamics package Wallerberger et al. (2019). The latter further serves as a reference system for the diagrammatic expansion that self-consistently incorporates non-local electronic correlations, including magnetic fluctuations. This approach allows the mutual influence of collective electronic fluctuations on single-particle quantities and vice versa Stepanov et al. (2022); Vandelli et al. (2024); Chatzieleftheriou et al. (2024); Stepanov et al. (2024b), yielding reliable results for both single- and two-particle observables in a broad regime of system’s parameters Harkov et al. (2021); Vandelli et al. (2022); Vandelli (2022).
@@
The second important outcome of our results is related to the overall behavior of the spin susceptibility across the BZ. As discussed above, according to experimental measurements the magnetic signal can be decomposed into the sum of the SDW peaks and a broad dome structure centered around the point, while DMFT calculations instead find a quasi-constant background signal besides the peaks. The D-TRILEX calculations reveal a significantly diminished spin susceptibility at the edges of the BZ, with a “cross”-like structure in momentum space of higher intensity appearing at the center of the zone, visible in the inset of Fig. 2. The overall structure of the susceptibility agrees very well with the experimental results Steffens et al. (2019). While early-unpolarized INS studies reported features of the spin excitation spectrum around Braden et al. (2002a), the most recent polarized INS data Steffens et al. (2019); Jenni et al. (2021) unambiguously confirmed the existence of such fluctuations, but cannot resolve the actual structure around . Nevertheless, strong spin excitations around are found in metamagnetic Ca2-xSrxRuO4 Friedt et al. (2004); Steffens et al. (2007, 2011), where the isostructural Ca/Sr substitution is, however, accompanied by a rotation and deformation of the RuO6 octahedra Friedt et al. (2001); Kriener et al. (2005).
@@
In Fig. 3 we show the orbitally resolved static spin susceptibility , where . The results are calculated using D-TRILEX (blue) and DMFT (green) and plotted along the X--X path. Although the incommensurate SDW peaks are present in all orbital components due to mixing of different of orbital contributions to the susceptibility, the value of is the largest for the intra-orbital - and - components (a). This is expected, since the SDW peaks originate from the nesting of the FS of these 1D-like orbitals. Interestingly, the background dome-shaped magnetic signal, found experimentally and reproduced by D-TRILEX, appears to originate from magnetic fluctuations within the 2D-like orbital (b). Instead, the DMFT results do not exhibit any significant signal around the point, and for all the components a quasi-local response is obtained besides the peaks. This result may have significant implications for the symmetry of the superconducting order parameter, as recent studies suggest that magnetic fluctuations within the orbital are a key ingredient for the electron pairing Rømer et al. (2019); Profe et al. (2024). In particular, the presence of spin excitations at more than one wave-vector allows for a competition between different order parameters of the superconducting state Ghosh et al. (2021); Benhabib et al. (2021).
@@
By analyzing the electron self-energy we find that, although it is not local and exhibits a noticeable momentum dependence perpendicularly to the FS, it has a surprisingly small momentum-dependence along the FS, as discussed in the SM SM . Along the FS the / orbitals appear to have a practically constant self-energy, while the orbital exhibits a finite but small momentum-modulation. This result is in a qualitative agreement with recent ARPES data presented in Fig. 4 (d) of Ref. Tamai et al. (2019), and explains why DMFT was particularly successful in reproducing the FS of the material.
@@
These excitations are found to be significant in this material, such that DMFT calculations predict an ordered SDW state that is not observed experimentally.
We demonstrate that the self-consistent inclusion of spatial magnetic fluctuations suppresses their strength by reducing many-body correlations to the extent that no magnetic ordering is realized in Sr2RuO4.
The overall behavior of the spin susceptibility in momentum space deduced from INS measurements is well reproduced by our calculations.
@@
We identify the orbital character of the unusual dome structure as resulting predominantly from the 2D-like -orbital. This observation resonates with recent claims that it is the -orbital that
is mainly responsible for the superconductivity.